Refractive index of water with respect to air is 1.33 and that of diamond is 2.42. (i) In which medium does the light move faster, water or diamond? (ii) What is the refractive index of diamond with respect to water?

Grade:10

1 Answers

Gaurav Yadav
3 years ago
Hey student.
 
Refractive index = speed of light in vacuum / speed of light in medium. Since the refractive index of diamond is more, hence the speed of light is lesser in diamond. Let speed of light in water vevw and in diamond bevd. Refractive index of diamond w.r.t water is say n = Speed of light in water / speed of light in diamond. n = vw/vd Dividing both numerator and denominator by speed of light [c] we get n = (vw/c) ÷ (vd/c) = Inverse Ratio of refractive index of water and diamond. n= 2.42/1.33 = 1.82 (approx.)
 
This is the answer to your question. Hope this helps you.